OpenStudy (anonymous):

Graph the six terms of a finite series where a1 = −3 and r = 1.5.

OpenStudy (anonymous):

@mathstudent55

OpenStudy (welshfella):

sum of n terms = a1 (r^n - 1) ------ r - 1

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

-3.5(1.5-1)/1.5-1 ?

OpenStudy (welshfella):

-3 * (1.5^6 - 1) ---------- 1.5 - 1

OpenStudy (anonymous):

I got 62.34???

OpenStudy (welshfella):

oh thats the sum of 6 terms sorry I misread the question

OpenStudy (anonymous):

It's fine :)

OpenStudy (welshfella):

its -62.34

OpenStudy (welshfella):

the terms in sequence are -3 , -3*1.5 , -3*1.5^2 , -3*1.5^3 ... = -3 , -4.5 , -6.75 , - 10.125 ....

OpenStudy (anonymous):

I see, so then I would just look for a graph that showed that?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

i have a feeling I'm wrong haha! So I'm going to go with D

OpenStudy (welshfella):

D is correct

OpenStudy (welshfella):

first 3 points are (1,-3), (2, -7.5) , (3, -14.25

OpenStudy (welshfella):

the y values are accumulated

OpenStudy (welshfella):

- so , for example, the second y value is -3 - 4.5 = -7.5

OpenStudy (anonymous):

Ohh! I understand now:) Thanks so much for your help!

OpenStudy (welshfella):

yw